### v1

# Super-intelligence or Superstition? Exploring Psychological Factors Underlying Unwarranted Belief in AI Predictions

Abstract
This study investigates psychological factors influencing belief in AI predictions about personal behavior, comparing it to belief in astrology and personality-based predictions. Through an experiment with 238 participants, we examined how cognitive style, paranormal beliefs, AI attitudes, personality traits, and other factors affect perceived validity, reliability, usefulness, and personalization of predictions from different sources. Our findings reveal that belief in AI predictions is positively correlated with belief in predictions based on astrology and personality psychology. Notably, paranormal beliefs and positive AI attitudes significantly increased perceived validity, reliability, usefulness, and personalization of AI predictions. Conscientiousness was negatively correlated with belief in predictions across all sources, and interest in the prediction topic increased believability across predictions. Surprisingly, cognitive style did not significantly influence belief in predictions. These results highlight the "rational superstition" phenomenon in AI, where belief is driven more by mental heuristics and intuition than critical evaluation. We discuss implications for designing AI systems and communication strategies that foster appropriate trust and skepticism. This research contributes to our understanding of the psychology of human-AI interaction and offers insights for the design and deployment of AI systems.
Keywords AI  $\cdot$ Large Language Model  $\cdot$ Cognitive Biases  $\cdot$ Prediction  $\cdot$ Belief
1 Introduction
Technology is often associated with scientific advancement and frequently viewed in opposition to superstition. However, the recent rapid advancements in artificial intelligence (AI) have given rise to quasi-religious perspectives among some leaders in the tech industry [1]. In fact, as generative AI and large language models (LLMs) continue to evolve, a contemporary critique posits that the narrative surrounding their development bears striking resemblances to religious discourse [2]. Promoting concepts such as superintelligence and sentience, prominent figures in the tech industry have been observed making prophetic claims regarding AI’s potential to either salvage or obliterate humanity [3, 4]. These claims are often based on personal speculation and subjective views rather than rigorous scientific research. Nonetheless, empirical research has demonstrated that these popular narratives are mirrored in public perceptions of AI, frequently manifesting as exaggerated utopian or dystopian visions [5, 6]. This intersection of technological advancement and quasi-religious perspective presents a complex landscape for scholars to navigate, as it blurs the boundaries between scientific progress and speculation.
While AI has been developed through predictive algorithms and statistical pattern recognition, exemplifying advancements in mathematical and rational thinking, public perception of AI may draw unexpected parallels with astrology [7, 8, 9]. Despite astrology representing an opposing epistemology based on pseudoscientific methodologies antithetical to empirical reasoning, both AI and astrology function as prediction-generating mechanisms helping people make sense of the uncertainty in the world [7, 8]. Anthropologist Christophe Lazaro suggests that AI, powered by data, has become a new mode of speculative thinking, with algorithms taking on the role of oracles in contemporary societies, framing AI as "artificial divination" [8, p. 128].
On the other hand, AI, unlike astrology, presents an aura of scientific legitimacy, backed by a vast body of scientific research in computer science, mathematics, cognitive science, and more. This foundation contributes to the belief that AI systems are rational and objective. However, studies have shown that algorithms could produce unpredictable outcomes, biased results, and inherit implicit societal biases [10, 11, 12]. Moreover, despite ongoing efforts to increase transparency, much of the inner workings of AI systems remain "black boxes" [13, 14, 15, 16]. Such misconceptions regarding AI may engender a form of naive technological optimism and irrational trust in the system, which Wilson refers to as "rational superstition" [17]. Scholars have investigated the concept of algorithmic appreciation—the propensity to rely on algorithms for decision-making processes—which may lead to an overreliance on AI in critical decision-making contexts [18, 19].
Popular narratives that over-inflate the promises of AI and its capabilities is particularly alarming when considering the impact it has on mental models and perception of AI systems [6]. According to the dual process theory in psychology, people engage in two types of thinking—fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Incomplete or incorrect mental models shaped by speculative claims about AI rather than scientific evidence can become a mental heuristic that triggers people’s System 1 thinking rather than engaging their System 2 thinking, blurring the lines between speculation and informed reasoning [22]. This, in turn, can lead to miscalibration of trust and over-reliance on AI, which drives ineffective use of AI [18, 19, 23].
For instance, consider an AI system that claims to give you personalized product recommendations based on your social media usage and purchase history. A user with an incorrect mental model of the system’s capabilities and limitations may hastily rely on its recommendations without considering other options. On the flip side, a user who has developed an overly skeptical mental model of AI may reject helpful recommendations and miss critical opportunities. Therefore, without ongoing effort to establish appropriate mental models, users can become more prone to manipulation and potential negative outcomes.
While belief in AI supremacy is growing despite its demonstrated limitations and biases, our study presents the first empirical investigation into the phenomenon of "rational superstition" [17] in AI. To do so, we examine how people’s perception and trust in AI predictions may correlate with other irrational behaviors, such as belief in astrology. This research question is particularly timely and relevant given the recent surge in popularity of astrology, especially among Millennials and Gen Z [24, 25], despite the lack of empirical evidence supporting its validity [26, 27]. A recent consumer survey of over 2000 people in the US revealed that about a third of Millennials and a quarter of Gen Z have made financial decisions based on their horoscope [28].
To address this research question, we conducted an experiment involving 238 participants who were presented with fictitious predictions from three sources (AI, astrology, and personality psychology) and rated their perceived validity, reliability, usefulness, and personalization. To create a reasonable basis for the predictions, participants answered questionnaires about astrology and personality and engaged in a simulated investment game that appeared to analyze their interactions and predict their future investment behavior. Participants were randomly assigned to either positive (N = 119) or negative (N = 119) prediction groups, receiving forecasts about their future investment behavior and outcomes accordingly.
Our study also explores how cognitive style, paranormal beliefs, gullibility, trust in AI, and personality traits may influence belief in predictions based on AI, astrology, and personality psychology. This investigation builds upon previous work examining how user characteristics such as personality traits influence the perception of AI-generated advice [29] and trust in AI [30]. Additionally, we consider the relationship between cognitive style and paranormal beliefs, including astrology [31, 32, 33, 34, 35], as well as how cognitive reflection affects social media usage and the sharing of (mis)information and conspiracy theories [36, 37].
Furthermore, we explore how demographic factors such as age, gender, and education level, as well as prior experience in AI, astrology, and personality psychology, and the level of interest in the topic influence belief in the predictions. By examining these various factors, our research aims to provide a comprehensive understanding of the "rational superstition" phenomenon in AI and its implications for decision-making in an increasingly AI-driven world.
Based on our research questions, we tested the following hypotheses:
  • H1: Individuals who are more likely to believe in predictions based on astrology and personality will be more likely to believe in AI predictions.
  • H2: Individuals with higher cognitive reflection and need for cognition will find AI predictions more credible and reliable than astrology- or personality-based predictions.
  • H3: Demographic factors and personal attributes (cognitive reflection, trust in AI, paranormal beliefs, gullibility, personality, level of interest, and prior experience) will moderate the perceived validity, reliability, usefulness, and personalization of predictions from different sources.
The study’s findings contribute valuable insights to the ongoing discourse on AI ethics, transparency, and responsible implementation. By shedding light on the factors that influence trust and belief in AI predictions, we aim to inform policymakers, AI developers, and educators about potential pitfalls and areas for improvement in AI literacy and public engagement with AI technologies.
2 Results
2.1 People who are more likely to believe in astrology and personality-based predictions are more likely to believe in AI predictions.
To test the first hypothesis (H1), a multiple linear regression model was applied to the data in wide format, with believability score for AI predictions as the outcome variable and believability scores for astrology- and personality-based predictions as predictor variables, along with control variables (cognitive style, paranormal beliefs, gullibility, AI attitude/trust in AI, big five personality, familiarity with the prediction sources (AI, astrology, personality), interest in topic of prediction, age, gender, education level). The scores across the four subscales (perceived validity, reliability, usefulness, and personalization) were averaged as overall "believability" scores to provide a straightforward test of the hypothesis.
The multiple linear regression analysis explained a significant proportion of the variance in the AI overall score ($R^{2}=0.7606$, Adjusted $R^{2}=0.7337$, $F(24,213)=28.2$, $p<0.001$). Significant predictors included the zodiac overall score (Estimate = 0.3119, $p<0.001$) and the personality overall score (Estimate = 0.4585, $p<0.001$), supporting the hypothesis that belief in astrology and personality-based predictions is positively associated with belief in AI predictions. This association can be visually observed in the scatterplots in Figure 1.
Figure 1: Relationship between belief in AI predictions and astrology predictions (left) and belief in AI predictions and personality predictions (right). The scatterplots show individual scores, with a linear regression line indicating a positive correlation. The shaded area represents the 95% confidence interval for the regression line.
2.2 People generally find fictitious AI predictions about their personal behavior convincing.
To examine the influence of moderating factors including cognitive style (H2), paranormal beliefs, trust in AI, personality traits, demographic factors, etc. (H3), a mixed-effects model was fitted to the data in a nested long format, with prediction source (AI, astrology, personality) and believability subscale (perceived validity, reliability, usefulness, and personalization) as categorical predictors, and subscale score as the outcome variable. This approach allowed examination of the main effects and moderating effects of psychological and cognitive factors on belief in AI predictions.
The model included fixed effects for subscale, prophecy source, prophecy group, cognitive style, paranormal beliefs, gullibility, AI attitude/trust in AI, big five personality, familiarity with sources, interest in topic, age, gender, education level. Interaction terms were included to assess the combined effect of prophecy source and select moderating factors on subscale scores. The model also included random effects to account for individual variability, with the random intercepts capturing the baseline differences among subjects and the random slopes capturing the variability within each participant across different conditions. The inclusion of random intercepts and slopes improved the model fit. For more on our methodology, see Section 7.
Based on the mixed effects model analysis, the baseline subscale score, associated with "AI" prophecy source, "Validity" subscale, "Positive" prophecy group, "Female" gender, and "Bachelor" education level, was 5.10 ($p<0.001$) on a 7-point Likert scale. The main effects of prophecy source on believability of AI predictions showed that on average, perceived validity of astrology-based predictions was 0.66 points lower compared to AI predictions (95% CI [-0.93, -0.39], p < 0.001). The difference was not statistically significant for personality-based predictions (0.07, 95% CI [-0.18, 0.32], p = 0.593). These results suggest that people generally found fictitious predictions from AI, astrology, and personality convincing.
There were significant main effects of subscales on the perception of AI predictions. On average, perceived personalization was rated 0.23 point higher than perceived validity (95% CI [0.08, 0.38], p = 0.002). On the other hand, perceived reliability was rated 0.91 points lower than perceived validity (95% CI [-1.13, -0.68], p < 0.001), and perceived usefulness was rated 0.63 points lower than perceived validity (95% CI [-0.83, -0.44], p < 0.001). This relationship between the subscales by prediction source can be visualized in Figure 2. The interaction between subscales and prophecy source showed that the main effects of subscales were largely consistent across AI, astrology, and personality, with some subscales being more pronounced for astrology. Perceived personalization was further increased by 0.23 points (95% CI [0.02, 0.44], p = 0.035) and perceived reliability further decreased by 0.31 (95% CI [-0.62, -0.00], p = 0.048) for astrology-based predictions, compared to the AI baseline. Other interaction terms were not significant, suggesting that the effects are not significantly different from the main effects.
Furthermore, we found that compared to the baseline “Positive” prediction group, the “Negative” prediction group was associated with a decrease in perceived validity by 1.19 points (95% CI [-1.53, -0.85], p < 0.001). We found no significant interactions between prophecy group and prophecy source, suggesting that this effect was consistent across AI, astrology, and personality. This suggests that while people generally found fictitious predictions believable, it was not the case for those who received negative predictions.
Figure 2: Boxplot showing the distribution of subscale scores by prophecy source and subscale. The boxes represent the interquartile range, with the means indicated in green triangles and notches representing the 95% confidence intervals for the medians.
2.3 There is no evidence of correlation between belief in predictions and cognitive style.
Based on our hypothesis (H2), we expected to see either a significant positive association between cognitive style and belief in AI predictions, significant negative interactions between cognitive style and other prediction sources (astrology, personality), or both. However, our results showed that the composite cognitive score, as measured as the composite of the performance-based Cognitive Reflection Test (CRT-2) [38] and the preference-based Need for Cognition (NFC-6) [39], did not significantly increase the perceived validity of AI predictions. The estimated increase in perceived validity was 0.13 points with one point increase in the composite cognitive score, but this effect was not statistically significant (95% CI [-0.01, 0.26], p = 0.065). The composite cognitive score ranged from -4.29 to 2.61 (Mean = 0.0, SD = 1.44).
However, we found some significant negative interactions between cognitive score and the subscales. Compared to perceived validity, higher composite cognitive score was associated with a decrease in perceived reliability (-0.11, 95% CI [-0.20, -0.02], p = 0.021) and perceived usefulness (-0.12, 95% CI [-0.20, -0.04], p = 0.004), making the main effect less positive. The interaction with perceived personalization was not statistically significant (-0.05, 95% CI [-0.11, 0.01], p = 0.092).
The interactions between cognitive style and prediction sources showed that while there was some negative interaction for astrology (-0.10, 95% CI [-0.21, 0.01], p = 0.079) and personality (-0.07, 95% CI [-0.18, 0.03], p = 0.170), the effects were not statistically significant. This suggests a lack of evidence that cognitive style is an influential factor in how people perceive predictions based on AI, astrology, and personality. Three-way interactions between subscales, prophecy source, and composite cognitive score did not lead to significant results.
2.4 Higher paranormal beliefs increases perceived validity, reliability, usefulness, and personalization of AI predictions.
One of the most interesting findings was that having paranormal beliefs was positively associated with belief in AI predictions. The results from the mixed effects model showed that the paranormal beliefs score, as measured by a shortened version of the Revised Paranormal Belief Scale (R-PBS) [40], significantly increased perceived validity of AI predictions. With each point increase in the paranormal beliefs scale, perceived validity of AI predictions increased by an average of 0.02 points (95% CI [0.01, 0.03], p = 0.001) on a 7-point scale. The paranormal score, which was centered for the analysis, originally ranged from 15 to 95 (Mean = 45.6, SD = 20.08).
The interaction between paranormal beliefs and subscales showed that the effect was stronger for perceived reliability (0.01, 95% CI [0.00, 0.01], p = 0.020) and perceived usefulness (0.01, 95% CI [0.01, 0.02], p < 0.001), both being statistically significant. There was no significant interaction effect for perceived personalization (0.00, 95% CI [-0.00, 0.01], p = 0.427), suggesting that the effect did not differ significantly from perceived validity.
Paranormal beliefs score was an even stronger predictor for belief in astrology-based predictions, perhaps less surprisingly given that the scale includes questions around astrology (see Section 7.3). Each standard deviation increase in paranormal beliefs led to an additional 0.01 point increase in perceived validity (95% CI [0.01, 0.02], p = 0.001) compared to the AI baseline. Differences for personality-based predictions were not significant compared to the AI prediction baseline (-0.00, 95% CI [-0.01, 0.00], p = 0.274). No significant interactions were observed between the subscales, prophecy source, and paranormal score.
2.5 Positive AI attitudes increases belief in AI predictions, especially perceived reliability.
Individuals with more positive attitudes towards AI found predictions based on AI more believable. One point increase in AIAS score [41] led to an increase of perceived validity of AI predictions by 0.04 points (95% CI [0.01, 0.06], p = 0.001) on a 7-point scale. The AIAS score ranged from 4 to 40 (Mean = 25.79, SD = 8.68).
The main effect was augmented by an additional 0.03 points for perceived reliability (95% CI [0.02, 0.05], p < 0.001), while the interactions were not significant for perceived personalization (-0.00, 95% CI [-0.01, 0.01], p = 0.941) and usefulness (0.01, 95% CI [-0.00, 0.03], p = 0.071). This suggests that positive attitudes toward AI/trust in AI is more closely related to perceived reliability among all subscales, which is aligned with the findings of studies that explore the relationship between trust/attitude and reliance in AI [42, 43, 44, 45, 46].
The interaction between AI attitude (AIAS) and prediction source showed that the positive impact of positive AI attitudes on perceived validity was mostly reversed for astrology-based predictions (-0.24, 95% CI [-0.41, -0.08], p = 0.004), while the effect did not differ significantly for personality-based predictions (0.04, 95% CI [-0.11, 0.19], p = 0.590). The interaction effects between subscales, prophecy source, and AI attitude score were not statistically significant.
2.6 People with high conscientiousness are less likely to believe in predictions about personal behavior.
Out of the big five personality traits (Extraversion, Openness, Agreeableness, Conscientiousness, Emotional stability), conscientiousness was negatively associated with perceived validity of predictions across all sources. With each point increase in the conscientiousness score (Mean = 5.29, SD = 1.32, Range = [1.0, 7.0], perceived validity was estimated to decrease by an average of 0.15 points (95% CI [-0.30, -0.01], p = 0.032) on a 7-point scale. There were no significant interaction effects with subscales observed, suggesting the main effect was consistent across the four subscales.
While the other domains of the five-factor model were not found to have significant influence, there were some variation in interaction terms. Extraversion was associated with a 0.06 point increase in perceived usefulness (95% CI [0.02, 0.10], p = 0.004), while Openness was associated with a 0.06 decrease in perceived personalization (95% CI [-0.10, -0.01], p = 0.012).
2.7 Level of interest in the prediction topic increases perceived validity, reliability, usefulness, and personalization.
While often overlooked in the context of studying trust in AI predictions, individuals’ interest in the topic of prediction (in our case, personal investing behavior) was an influential factor in perceived validity, personalization, reliabiliity, and usefulness of predictions. With each point increase in the level of interest in the topic, perceived validity significantly increased on average by 0.27 points (95% CI [0.11, 0.43], p = 0.001) on a 7-point scale. The interest in behavior scale ranged from 1 to 5 (Mean = 3.18, SD = 1.09).
The positive association was observed across other subscales. Compared to perceived validity, perceived personalization increased slightly less, reduced by 0.06 points (95% CI [-0.11, -0.01], p = 0.018), while interaction was not statistically significant for perceived reliabilty (-0.07, 95% CI [-0.14, 0.01], p = 0.068) and perceived usefulness (0.02, 95% CI [-0.04, 0.09], p = 0.475), suggesting that the effects were comparable to perceived validity. The results suggest that when all else is held constant, the more people are interested in the topic, the more likely they will perceive fictitious predictions to be valid, reliable, useful, and personalized.
2.8 There is no evidence that the level of familiarity with the prediction sources influences belief in predictions.
On the other hand, familiarity with the prediction sources (AI, astrology, personality psychology), which measures the level of self-reported familiarity/prior knowledge in each source, did not have a significant effect on perceived validity (-0.02, 95% CI [-0.11, 0.07], p = 0.650) across all sources. The familiarity scale ranged from 1 to 5 (Mean = 2.96, SD = 1.05).
The interaction between familiarity and subscales shows that perceived personalization increases slightly but significantly by 0.05 points (95% CI [0.00, 0.11], p = 0.035), while differences were not statistically significant for perceived reliability (0.02, 95% CI [-0.06, 0.10], p = 0.685) and perceived usefulness (0.01, 95% CI [-0.06, 0.08], p = 0.781).
Unlike what the literature suggests about the familiarity effect, a cognitive phenomenon where people tend to prefer things they are familiar with [47], our results showed inconclusive evidence that familiarity in the prediction sources (AI, astrology, and personality) neither increased nor decreased belief in the respective predictions.
2.9 Other results
Gullibility was not found to be a significant predictor of perceived validity of AI predictions (-0.02, 95% CI [-0.04, 0.01], p = 0.128). While the interaction between gullibility and Personalization subscale was significant (0.01, 95% CI [0.00, 0.02], p = 0.021), the results were inconclusive when compared to the main effects. Interactions between gullibility and prophecy source were not significant (Astrology: -0.00, 95% CI [-0.02, 0.02], p = 0.977, Personality: -0.02, 95% CI [-0.04, 0.00], p = 0.096), suggesting the effects were similarly insignificant across different prophecy sources. Gullibility scale ranged from 6 to 39 (Mean = 14.68, SD = 7.69).
We found that older age was associated with a decrease in belief in predictions across all sources. One year increase in age (Mean = 41.04, SD = 12.38, Range = [19, 75]) was associated with a decrease of perceived validity of AI predictions by 0.02 points (95% CI [-0.03, -0.00], p = 0.014). Interactions between age and subscales were not significant, nor were interactions between age and prophecy source. The non-significant interaction terms suggested that the effect was consistent across subscale and prophecy source.
For gender, we did not find significant main effects relative to the Female reference level. While male participants were associated with lower perceived validity scores for AI predictions than female participants, it was not statistically significant (-0.31, 95% CI [-0.69, 0.06], p = 0.103). However, they tend to perceive them as more reliable than female participants (0.41, 95% CI [0.14, 0.68], p = 0.003). Interactions with other subscales were not significant, suggesting they were similar to the perceived validity baseline. Interactions between gender and prophecy source were not significant.
There were no significant main effects for level of education compared to the Bachelor reference level. Interactions between education level and subscale showed some significant effects between High school and Personalization (-0.23, 95% CI [-0.40, -0.06], p = 0.009), and Less than high school and Reliability (-0.89, 95% CI [-1.56, -0.22], p = 0.009), but these results did not lead to conclusive results.
3 Discussion
Belief is subjective and context-dependent; it is also socially constructed through external stimuli such as media narratives, contextual factors, and internal disposition [48, 49]. When looking at belief, trust, and reliance in AI predictions, it is important to consider not just the features of the AI system, but also the characteristics of the users and the broader context. The results from our study show that people who are more likely to believe in astrology- and personality-based predictions were more likely to believe in AI predictions. Moreover, we found that belief in AI predictions on personal behavior can be influenced by various psychological, cognitive, and contextual factors that also drive belief in astrology- or personality-based predictions. By breaking down the concept of believability into four subscales (perceived validity, reliability, usefulness, and personalization), which we further elaborate in Section 7.3.1, we were able to observe interesting patterns and interactions with different subscales.
Cognitive style. While prior literature seemed to suggest that an analytic cognitive style would lead to more skepticism in predictions based on astrology and possibly personality [31, 33, 35], our results did not show significant evidence to support these claims. Moreover, cognitive style was not a significant predictor in belief in AI predictions. Thus, our hypothesis (H2) that people with more analytic cognitive style would be more likely to believe in AI predictions than predictions based on astrology and personality was not supported by our findings. This suggests that analytic cognitive style may not necessarily lead to rational skepticism of fictitious predictions, and perhaps there may be a missing link or mechanism between cognitive style and belief/trust in AI predictions, which could be a topic of future study.
Paranormal beliefs. Paranormal beliefs were found to be an influential predictor of belief in predictions across all sources, including AI predictions. While the association may seem unlikely at first, it can be interpreted within the context of the influence of popular AI narratives. The popular portrayal of AI as rational and objective creates the impression that people would also treat AI in a rational way, and thus more scientifically-inclined people would find AI predictions more believable. However, our results show that belief in AI predictions is more closely associated with paranormal beliefs than one’s cognitive style. This result points to the existence of the phenomenon of "rational superstition" in AI that was described in the introduction. Moreover, while the effect of paranormal beliefs on belief in AI predictions was significant across all subscales, the effect was larger for perceived usefulness and reliability. Social science research suggests that the rise in belief in astrology is an indicator of rising uncertainties and anxieties as a result of disintegration of community in the modern era [50]. As such, people turn to pseudo-scientific approaches such as astrology for insights to help them navigate their lives. Our findings suggest that people may perceive AI predictions in a way that is similar to the way they perceive astrology-based predictions, which could help explain the results.
AI Attitude/Trust in AI. Our results showed that people with positive attitudes towards AI found AI predictions more valid, reliable, useful, and personalized. Among the subscales, perceived reliability was found to be most closely related to attitude towards AI. This supports findings from prior literature that positive attitudes toward AI lead to higher reliance in AI predictions [51, 46].
Personality. Our findings using the Big Five personality traits [52] show that conscientiousness had a negative influence on perception of validity, personalization, reliability, and usefulness, while other traits did not have significant effects. This contrasts previous findings by Riedl [30], who found that agreeableness, openness, and extraversion positively influenced trust and high neuroticism negatively impacted trust, while results on conscientious were mixed.
Conscientiousness is defined as “socially prescribed impulse control that facilitates task- and goal-directed behavior, such as thinking before acting, delaying gratification, following norms and rules and planning, organizing, and prioritizing tasks” [53, p. 120]. Prior studies have found a positive relationship between conscientiousness and cognitive ability [54, 55]. Contrary to our expectations that these two traits may mirror each other in direction, it was not the case—cognitive style did not have significant effect on believability in our study, while conscientiousness had a significant negative influence on perception of validity, personalization, reliability, and usefulness.
Interest in the topic of prediction. Our findings show that people’s interest in the topic of prediction is a strong predictor when it comes to their belief in predictions, regardless of the source. While our results do not confirm any causal relationship, it could suggest that the more interested one is in a topic, the more exposure they may have to information about it, and the more likely they may be susceptible to cognitive biases that could influence their belief. Some biases that could explain this tendency are confirmation bias and belief bias. Confirmation bias is when individuals seek out and remember information that confirm their existing beliefs [56]. Belief bias occurs when individuals judge the strength of an argument based on the believability of its conclusion rather than the logical structure of the argument itself [57, 58].
This has implications in the context of AI-assisted decision-making in different fields; e.g. in marketing, companies may take advantage of people’s interest in a certain topic and expose them to more targeted information (e.g. advertisements) that may create an impression of hyper-personalization and lead to more positive attitudes toward the brand [59]. Moreover, in the medical field, an individual may be more prone to believing AI-based predictions about their health if it confirms their prior beliefs.
Familiarity/Prior knowledge. Familiarity with the prediction source was not found to be a significant predictor of belief in predictions. Earlier, we discussed how people’s positive AI attitudes and interest in the topic of the prediction were positively correlated with belief in the predictions, which may be explained by its likelihood to fuel certain cognitive biases such as belief bias. On the other hand, an individual’s level of familiarity in AI, astrology, and personality psychology may not be necessarily indicative of their attitude towards the field. For instance, someone who is not acquainted with the details of latest AI developments may have either a utopian or dystopian view of AI’s impact on society, as mentioned in the introduction. Similarly, just because an individual is very familiar with astrology does not inherently mean they are more likely to believe or adopt the predictions, as that may depend on other personal and contextual factors. This suggests two things: 1) simply knowing more about the prediction source does not predict whether one finds a prediction more or less believable, and 2) it does not make one better at calibrating one’s trust, providing implications for the design of trustworthy AI systems.
Gullibility. While prior literature suggested a positive relationship between gullibility and paranormal beliefs [31, 34], our findings showed inconclusive results on whether self-reported gullibility had a positive or negative effect on the belief of fictitious predictions based on AI, astrology, and personality. There may be a few potential explanations for this. One is that the order of the study potentially biased participants’ answers. The survey was completed after participants had seen and evaluated the predictions, at which point their perception of the believability of the predictions may shadow their self-perception of gullibility—in other words, they did not want to contradict their decisions so quickly by admitting that they are more prone to being fooled. Another possible explanation is the limitation of a self-report nature of the gullibility scale. Some people may be hesitant to admit that they are gullible for different reasons. As such, future studies can further explore the relationship between gullibility and belief in AI predictions.
Demographic factors. Our results found that older people are more skeptical of predictions about personal behavior, across AI, astrology, and personality sources. This is aligned with prior literature that seems to suggest that older age is associated with lower trust in AI [60], due to many reasons including bias, lack of learning avenues, and concerns about privacy [61]. For gender, while the main effects were insignificant, interactions revealed that male participants were more likely to perceive AI predictions as reliable than female participants. This is supported by a prior study that found that being male with higher education and with a Western background were predictors of trust in AI among the general population [62]. Moreover, while prior work suggested a positive association between level of education and belief in AI [60], we did not find significant main effects for level of education, and the interaction terms did not lead to conclusive results.
4 General Discussion
Our results highlight the irrational side of how humans perceive and believe in AI predictions, by providing a side-by-side comparison of the perception of fictitious predictions based on AI, astrology, and personality. Our analysis showed that belief in AI prediction about personal behavior may be more related to paranormal beliefs than cognitive style, and that positive attitudes toward AI and interest in the topic of prediction, but not familiarity with the sources, can enhance that belief.
These findings suggest that highly accurate and reliable performance is not a prerequisite for people to put high trust in predictions; models do not need to be completely trustworthy for users to put trust in them. The disentanglement between perception and actual performance leads to a few important discussions regarding how people perceive and engage with AI systems:
1) Incorrect mental models of AI can override perception of validity.
Belief in AI predictions is based on people’s mental model of AI systems, rather than the actual model performance. Hoff and Bashir presented a comprehensive three-layer framework of trust in automation, including dispositional, situational, and learned trust, among which system performance is part of learned trust that comes with active engagement with the system [63]. In our study, we observed that even without any validation of system performance, people found fictitious AI predictions highly valid, reliable, useful, and personalized. Other researchers have studied this placebo effect of AI systems, with just priming with descriptions of AI leading to people perceiving and behaving differently [64, 65, 66, 67].
This highlights the importance of the role of mental models in how people perceive and trust AI [23, 68, 69, 70, 71, 72]. Mental models are the collection of beliefs and expectations around AI systems and interactions with them, and can be shaped by prior beliefs as well as ongoing interactions with AI systems [23, 68]. By framing the scenario in a certain way, including the task domain and the information provided about the AI system, we observed that people could perceive fictitious AI predictions to be highly accurate, personalized, and trustworthy [23].
Moreover, the results revealed that people might perceive AI predictions similarly to how they believe in astrology predictions, given a reasonable context. This supports the observation of "rational superstition" in AI and automation [17]. We argue that this phenomenon is fueled by an idealized portrayal of AI systems as enchanted yet rational and deterministic [73], which may shape people’s prior beliefs about AI and bias their interactions.
The dual process theory in psychology provides insight on how larger narratives could bias people’s perceptions of AI systems. The theory states people engage in two types of thinking—fast and intuitive (System 1) and slow and reflective (System 2) [20, 21]. Mental models may be shaped by evolving public narratives around AI [6], and may become a mental heuristic that triggers people’s System 1 thinking rather than engaging their System 2 thinking, blurring the lines between speculation and informed reasoning [22]. Incomplete or incorrect mental models shaped by speculative claims about AI rather than scientific evidence can lead to over-reliance or miscalibration of trust [18, 19], driving ineffective use of AI [23, 43].
This points to the importance of de-divinizing AI in popular narratives and increasing awareness of both the capabilities and limitations of AI systems to help users form more accurate mental models. Even with technological advancements, people’s mental models should evolve along with an appropriate level of skepticism to avoid becoming overreliant or overly skeptical.
2) Mental models can influence and be influenced by cognitive biases.
Cognitive biases can provide some insight into how people might believe in AI in an irrational manner. Specifically, belief bias occurs when individuals evaluate a statement based on its plausibility according to their prior beliefs rather than its logical validity [57, 58]. When situated within the dual process theory, belief bias is known to be largely attributed to System 1 thinking (especially under time pressure [74]), but it can also be intervened by System 2 thinking to construct mental models around existing beliefs in a way that can be biased, leading to overriding of logic [75, 21]. Known as the selective processing model [76], this theory explains why people may engage logic when a statement contradicts their beliefs, but fail to engage logical thinking when a statement aligns with what they believe, especially when it comes to beliefs about themselves.
Another cognitive bias that may explain our findings is the Barnum effect, which refers to when individuals believe vague and general statements to be highly accurate for them personally. Also known as the Forer effect or personal validation fallacy, this cognitive bias has been studied in contexts such as astrology and personality tests [77, 78]). Based on the findings from our study, we argue that the Barnum effect can also explain how people engage with AI predictions about oneself, where people perceive general AI-generated statements as personalized and valid. Cognitive biases such as belief bias and the Barnum effect can influence how mental models are formed and how users rely on AI systems [69]. On the other hand, incorrect mental models can also induce certain cognitive biases, as we have seen in our findings.
3) Design of AI systems must take cognitive biases and mental models into account to support appropriate level of trust.
Awareness of the potential impact of cognitive biases on human-AI interaction is crucial to reduce the risk of negatively exploiting these human tendencies. For instance, AI systems that simply validate users’ prior beliefs may be more likely to be favored and accepted, but this may not support better decision-making or well-being, such as in areas of personal finance or health. Thus, we could think about how to design systems that are not only agreeable, but can present information in a way that helps people think critically.
Moreover, establishing correct mental models about AI systems is important. For example, in the context of mental health chatbots, users with incorrect mental models may have unrealistic expectations about what the AI can do, thus leading to greater disappointments. AI systems should include mechanisms that help users develop an appropriate level of trust in these systems to avoid over or under reliance [46]. Fortunately, mental models are dynamic and can evolve with more interactions [23, 79].
There are also important considerations for designing explainable AI systems. AI explanations are considered as a means to augment predictions to establish appropriate level of trust. However, with inaccurate mental models, explanations can increase over-trust in AI systems, even if they are fake [80]. A prior study has shown that belief bias could explain why explanations of AI systems could lead to overreliance [81]. When explanations align with users’ prior beliefs, it is much more likely that they will engage System 1 without much critical thinking. Given this awareness, explanations can be designed in a way to help users engage critical thinking and establish an appropriate level of trust, such as the approach of questioning the user rather than simply providing a deterministic answer [82], or showing a set of examples that the AI model considers similar to the target and the relationships between the features [83].
4) Our understanding of human-AI interaction needs to include psychological and contextual factors.
How people believe, trust, and rely on AI predictions for their decision-making should be seen in a larger context [46]. Contextual factors, including individual differences, task domain, information provided, media discourse, organizational and social environment, and more are all factors that could influence user perceptions and behavior in AI-assisted decision-making [11, 23].
Going beyond system performance, our findings highlight the importance of user-related factors in the evaluation of human-AI interaction. Importantly, our study underscores the role of psychological factors that may influence perception and belief in AI predictions. We found that factors including prior attitude towards AI, tendency to believe in paranormal phenomena, certain personality traits, and the level of interest in the topic influence people’s perceptions of AI predictions. We discuss the findings in the context of broader societal narratives around AI, mental models, and cognitive biases.
5 Limitations and Future Work
While our study has presented valuable insights into the phenomenon of "irrational superstition" in AI predictions and cognitive biases, there are several important limitations that can be addressed in future work. First, our experiment was specifically designed to provide predictions on personal behavior. There are some limitations to the extent to which our results can be generalized to the general use of AI. Therefore, future studies could look at how people perceive AI predictions in different contexts and prediction topics, while also exploring the relationship with underlying psychological and cognitive factors at play. For instance, the effect can be tested within the context of AI-assisted decision-making in personal health or well-being, or in educational settings using an AI tutor. We believe that our study opens up more questions about understanding the intricate ways in which humans perceive, trust, and interact with AI systems.
Second, like most survey designs, our survey-based approach to measuring psychological and cognitive traits may be affected by social desirability bias (the tendency to respond in a way that seems socially favorable) [84], as well as common method biases [85], and reliance on self-reported data. Therefore, we hope future studies can validate our findings using different methods and approaches.
Third, we were unable to fully resolve the positive autocorrelation issue in our mixed effects model (see Section 7.5.3), which may pose a slight limitation to the interpretation of our results. Future studies could explore alternative approaches or additional data collection to better address this issue.
Lastly, while our study mainly focused on user-related factors, future studies could also look at other contextual factors such as social and organizational environments and their impact on user perception and trust.
6 Conclusion
Our study empirically investigated the phenomenon of “rational superstition” in AI by providing a side-by-side comparison of people’s perceptions of fictitious predictions based on AI, astrology, and personality. Even without any validation of system performance, people perceived fictitious AI predictions as highly believable, and their perceptions were positively associated with their perception of astrology- and personality-based predictions. Our findings showed that people’s belief in AI predictions was positively associated with paranormal beliefs, positive attitude towards AI, interest in the topic of prediction, and negatively associated with conscientiousness and age. We also explored interactions with other prediction sources and the subscales including perceived validity, reliability, usefulness, and personalization. In our discussion, we highlighted the role of mental models and cognitive biases in shaping people’s perceptions of AI predictions, and the importance of including psychological and contextual factors in studying human-AI interaction.
7 Methods
7.1 Participants
Data were initially collected from a survey administered to 300 participants recruited through Prolific. After excluding 62 participants due to incomplete responses or failed attention checks, the final analysis was conducted with data from the remaining 238 participants. These individuals were adults aged 18 and above from diverse socioeconomic backgrounds (see Table 1 for an overview of participants’ demographic information).
Number per condition Positive 119
Negative 119
Age Mean 41.04
SD 12.38
Gender Female 53%
Male 43%
Other 3%
Education High school 12%
Some college 21%
Bachelor 43%
Master 14%
Other 9%
Table 1: Summary of participants’ demographic information
7.2 Experiment Protocol
Figure 3: Dashboard of the simulated investment game
7.2.1 Assessments and simulated investment game
Participants were asked to complete a short zodiac-related questionnaire (date, hour, and location of birth), a short personality test (adaptation of Myers-Briggs Type Indicator (MBTI) [86], and a simulated investment game. In the instructions, participants were told that they will receive personalized predictions about their future investment behavior generated from three distinct sources (astrology, personality, and AI) based on their game interactions and the information they provided. What they were not told in the beginning was that these predictions were generic, pre-determined statements that were not based on their responses, which was revealed to them after completion.
The simulated investment game was designed to elicit interactions comparable to that of modern robo-advising platforms [87], but in a simplified way. Participants were provided with virtual currency of \$10,000 to invest across three investment categories (high risk/return, medium risk/return, low risk/return) over ten rounds (representing years). At each round, they could allocate up to 100% of their assets across the three categories, with the goal of maximizing their total portfolio value.
A few design choices were made to increase the sense of realism for the investing game and to increase engagement:
  • Recommended portfolio allocation: Participants filled out a short questionnaire about their risk preferences for investing before starting the game and received a "customized" portfolio recommendation that they could view throughout the game. In reality, all participants received the same portfolio recommendation.
  • Realistic market scenario: The game was designed so that participants could experience the hypothetical ups and downs of the market, through a pre-determined series of alternating "bull market" (more gains) and "bear market" (more losses) scenarios that influenced the probability of the possible investment outcomes at each round.
  • Market forecasts: To further enhance the emotional engagement, each round started off with either a positive or negative market forecast that were aligned with the behind-the-scenes market scenario. A detailed market scenario that was used in the game and examples of the market forecast can be found in the appendix [ref].
  • Financial incentives: To motivate engagement, participants were told their investment performance would increase their total compensation for the study, inspired by previous studies that used real-life incentives to increase engagement [88, 89].
7.2.2 Predictions
After the game, participants were presented with pre-determined, fictitious predictions from three different sources (AI, astrology, personality) in randomized order. Participants were randomly assigned to either the “Positive” prediction group (rational investor, higher returns) or “Negative” prediction group (impulsive investor, lower returns). Examples of predictions received are shown in Table 2. The initial predictions were created by the authors, with variations generated using ChatGPT-4. These variations were randomly presented to enhance variability and simulate a placebo effect of AI-generated outputs. For each prediction, they were asked to respond to evaluate its perceived validity, personalization, reliability, and usefulness. The evaluation questionnaire is provided in Section 7.3.1.
Source Positive prediction group Negative prediction group
Astrology As the Moon’s celestial path brings it into close proximity with the stoic Saturn, this cosmic alignment sheds light on the inherent human struggle between emotion and reason when making financial decisions. Based on your astrological sign, you are more likely to make rational investment decisions based on reason rather than emotion. You have a good grasp of your emotions especially when it comes to decisions about money. Therefore, we predict that your portfolio will have higher-than-average performance in the long run. As the Moon’s celestial path brings it into close proximity with the stoic Saturn, this cosmic alignment sheds light on the inherent human struggle between emotion and reason when making financial decisions. Based on your astrological sign, your investment decisions are often impulsive, swayed by emotions rather than thoughtful analysis. Your emotions play a significant role in your investment decisions. As a result, it’s expected that your portfolio’s performance will be under the average in the long run.
Personality Whether you’re an analytical thinker or led by intuition, your Myers-Briggs profile offers valuable insights into your behavior when it comes to making investment decisions. Based on your personality analysis, you are more likely to make rational investment decisions based on reason rather than emotion. You have a good grasp of your emotions especially when it comes to decisions about money. Therefore, we predict that your portfolio will have higher-than-average performance in the long run. Whether you’re an analytical thinker or led by intuition, your Myers-Briggs profile offers valuable insights into your behavior when it comes to making investment decisions. Based on your personality analysis, your investment decisions are often impulsive, swayed by emotions rather than thoughtful analysis. Your emotions play a significant role in your investment decisions. As a result, it’s expected that your portfolio’s performance will be under the average in the long run.
AI Within the digital echo of your actions, AI sifts through the noise to forecast your next move. It transforms your clicks, likes, and digital whispers into a map of future investment choices. Based on our AI model analysis, you are more likely to make rational investment decisions based on reason rather than emotion. You have a good grasp of your emotions especially when it comes to decisions about money. Therefore, we predict that your portfolio will have higher-than-average performance in the long run. Within the digital echo of your actions, AI sifts through the noise to forecast your next move. It transforms your clicks, likes, and digital whispers into a map of future investment choices. Based on our AI model analysis, your investment decisions are often impulsive, swayed by emotions rather than thoughtful analysis. Your emotions play a significant role in your investment decisions. As a result, it’s expected that your portfolio’s performance will be under the average in the long run.
Table 2: Example of predictions received by Positive and Negative prediction groups
7.2.3 Post-study questionnaire
Following the experiment, participants were asked to respond to a series of questionnaires on participant characteristics (cognitive style, paranormal beliefs, gullibility, trust in AI/attitudes towards AI, personality, etc.) and demographic information (age, gender, education level). A summary of the scales used in the questionnaires is provided in Table 3. Detailed questionnaires can be found in Section 7.3.
Variable name Measure
Cognitive Style A composite measure of cognitive style that combines cognitive reflection test (CRT-2) [38] and need for cognition (NCS-6) [39].
Paranormal Beliefs A measure of participants’ belief in paranormal phenomena. A shortened version of R-PBS [40] was used.
AI Attitude Score A score indicating participants’ attitudes towards AI predictions. AIAS-4 [41], a 4-item scale was used. Higher score indicates more positive attitudes.
Gullibility A self-reported measure of participants’ tendency to be gullible, developed by [90] .
Big Five Personality Traits Includes extraversion, openness, agreeableness, conscientiousness, and emotional stability. Used 10-item Big Five Personality Inventory (TIPI) [91].
Interest in behavior A measure of participants’ interest in the topic of prediction (personal investment behavior). "How would you rate your level of interest in understanding your future investment behavior?" (5-point Likert scale, 1=Not interested, 5=Extremely interested)
Familiarity A measure of participants’ familiarity with the prediction sources (AI, Astrology, Personality). "How would you rate your level of familiarity with the following: 1) Astrology and horoscopes, 2) Personality psychology, 3) AI prediction systems" (5-point Likert scale, 1=Not at all, 5=Extremely familiar)
Table 3: Summary of the participant characteristics and their measures
7.3 Measures
This section includes the detailed questionnaires that we used for the study.
7.3.1 Perceived validity, reliability, usefulness, and personalization
To measure the believability of the predictions, we presented a short questionnaire that assess perceived validity, reliabiliy, usefulness, and personalization of given predictions. These factors were chosen based on existing literature on the different qualities of a statement that contribute to persuasion, user adoption of an information technology, and positive attitudes.
Perceived validity/accuracy of the content and reliability of the source was added according to the Elaboration Likelihood Model (ELM) in sociology, which argues that the logical validity of an argument is considered central to its believability, but people can also engage in peripheral ways of being persuaded, such as the reliability of the source [92]. Moreover, perceived usefulness was included as a metric that contributes to believability of a prediction, inspired by the Technology Acceptance Model (TAM)[93]. According to this model, perceived usefulness, defined as the degree to which a person believes that using a particular system would enhance their job performance, significantly influences users’ attitudes and intentions, which in turn affect actual technology usage. Lastly, perceived personalization has been found to increase attention and positive attitudes toward a statement [94].
We asked participants to rate eight statements on 7-point Likert scales (1=Strongly disagree, 7=Strongly agree), which include two statements per subscale (perceived validity, personalization, reliability, and usefulness).
  • I find the prediction convincing. (validity_1)
  • I can identify with the prediction. (personalization_1)
  • The source of the prediction is reliable. (reliability_1)
  • I find the prediction helpful. (usefulness_1)
  • The prediction is accurate. (validity_2)
  • The prediction describes me very well. (personalization_2)
  • I trust the source of the prediction. (reliability_2)
  • The prediction is useful for making future decisions. (usefulness_2)
7.3.2 Cognitive Reflection (CRT-2) and Need for Cognition (NCS-6)
Cognitive reflection, or the tendency to suppress an intuitive but incorrect response in favor of a more reflective and correct one, was chosen as a moderating factor for belief in predictions to explore how it influences the acceptance and trust in predictions about personal behavior. This approach aims to determine if higher cognitive abilities correlate with more critical evaluation and discernment of AI predictions, as well as predictions based on astrology and personality.
The Cognitive Reflection Test (CRT) was designed to predict performance in normative decision-making [95]. In our study, we used CRT-2 [38], an alternate form of the original 6-item CRT, in order to address issues of familiarity and learning effects that could potentially skew results. The 4-item CRT-2 maintains the same objective of measuring cognitive reflection but includes different questions to ensure a fresh and unbiased assessment, with less focus on numeric abilities [38].
The Need for Cognition (NFC) test is a self-report measure of an individual’s tendency to engage in and enjoy effortful thinking [96]. We use the 6-item NFC assessment (NCS-6) developed by [39].
While the CRT measures objective performance, the NFC measures subjective preference, which have been found to be correlated [97]. To account for the conceptual similarity of these two measures, they were combined to create a composite measure under the assumption that they have equal contribution to the final score. To do so, we summed the z-scores of the results from the two tests to account for the differences in range and variability [98].
Cognitive Reflection Test (CRT-2) [38]:
  • If you’re running a race and you pass the person in second place, what place are you in?
  • A farmer had 15 sheep and all but 8 died. How many are left?
  • Emily’s father has three daughters. The first two are named April and May. What is the third daughter’s name?
  • How many cubic feet of dirt are there in a hole that is 3’ deep x 3’ wide x 3’ long?
Need for Cognition (NCS-6) [39]:
  • I would prefer complex to simple problems.
  • I like to have the responsibility of handling a situation that requires a lot of thinking.
  • Thinking is not my idea of fun. (R)
  • I would rather do something that requires little thought than something that is sure to challenge my thinking abilities. (R)
  • I really enjoy a task that involves coming up with new solutions to problems.
  • I would prefer a task that is intellectual, difficult, and important to one that is somewhat important but does not require much thought.
7.3.3 Paranormal beliefs (R-PBS)
Paranormal beliefs refer to the conviction in phenomena beyond scientific explanation, such as supernatural events, astrology, and mystical experiences. Paranormal beliefs were chosen as a moderating factor for belief in AI, astrology, and personality predictions to investigate how these unconventional beliefs influence the acceptance of various predictive statements.
To quantify the degree of paranormal belief, we adapted the Revised Paranormal Belief Scale (R-PBS) [40] (7-point Likert scale, 1=Strongly disagree, 7=Strongly agree) to focus on the more relevant factors for our study, including traditional religious beliefs (4), spiritualism (4), precognition (4), superstition (3).
  • The soul continues to exist though the body may die. [religious beliefs]
  • Black cats can bring bad luck. [superstition]
  • Your mind or soul can leave your body and travel (astral projection). [spiritualism]
  • Astrology is a way to accurately predict the future. [precognition]
  • There is a devil. [religious beliefs]
  • If you break a mirror, you will have bad luck. [superstition]
  • During altered states, such as sleep or trances, the spirit can leave the body. [spiritualism]
  • The horoscope accurately tells a person’s future. [precognition]
  • I believe in God. [religious beliefs]
  • The number “13” is unlucky. [superstition]
  • Reincarnation does occur. [spiritualism]
  • Some psychics can accurately predict the future. [precognition]
  • There is a heaven and a hell. [religious beliefs]
  • It is possible to communicate with the dead. [spiritualism]
  • Some people have an unexplained ability to predict the future. [precognition]
7.3.4 AI Attitude/Trust in AI (AIAS-4)
A 4-item, 1-factor AI attitude scale (AIAS-4) [41] (10-point Likert scale, 1=Not at all, 10=Completely agree) was used to measure participants’ attitude towards AI.
  • I believe that AI will improve my life.
  • I believe that AI will improve my work.
  • I think I will use AI technology in the future.
  • I think AI technology is positive for humanity.
7.3.5 Gullibility
A shortened version of the originally 12-item, 2-factor scale by Teunisse et al. (2020) [90] (7-point Likert scale, 1=Strongly disagree, 7=Strongly agree) was used to measure self-reported gullibility.
  • (G4) I’m not that good at reading the signs that someone is trying to manipulate me. [insensitivity]
  • (G5) I’m pretty poor at working out if someone is tricking me. [insensitivity]
  • (G6) It usually takes me a while to “catch on” when someone is deceiving me. [insensitivity]
  • (G1) I guess I am more gullible than the average person. [persuadability]
  • (G9) My friends think I’m easily fooled. [persuadability]
  • (G12) Overall, I’m pretty easily manipulated. [persuadability]
7.3.6 Big Five Personality
Based on the Five Factor personality model, the 10-item Big Five Personality Inventory (TIPI) [91] (7-point Likert scale, 1=Strongly disagree, 7=Strongly agree) was used to measure participants’ personality.
I see myself as:
  • Extraverted, enthusiastic.
  • Critical, quarrelsome.
  • Dependable, self-disciplined.
  • Anxious, easily upset.
  • Open to new experiences, complex.
  • Reserved, quiet.
  • Sympathetic, warm.
  • Disorganized, careless.
  • Calm, emotionally stable.
  • Conventional, uncreative.
7.3.7 Familiarity/Level of expertise
How would you rate your level of familiarity with the following (5-point Likert scale, 1=Not at all, 5=Extremely familiar)
  • Astrology and horoscopes
  • Personality psychology
  • AI prediction systems
7.3.8 Interest in topic of prediction
How would you rate your level of interest in understanding your future investment behavior? (5-point Likert scale, 1=Not interested, 5=Extremely interested)
7.4 Approvals
This research was reviewed and approved by the MIT Committee on the Use of Humans as Experimental Subjects, protocol number E-5768.
7.5 Analysis
7.5.1 Variables
The dependent variable (Y) was the subscale scores for perceived validity, personalization, reliability, and usefulness (7-point Likert scale, 1=Strongly disagree, 7=Strongly agree).
The analysis included a range of predictor variables (X) to examine their effects on the subscale scores. The predictor and control variables were:
  • prophecy_source: A categorical variable indicating the source of the prophecy (Astrology, Personality).
  • prophecy_group: A categorical variable indicating the valence of prophecy (Positive, Negative).
  • composite_score: A composite measure of cognitive style.
  • paranormal_score: A measure of participants’ belief in paranormal phenomena.
  • aias_score: A score indicating participants’ attitudes/trust towards AI predictions.
  • gullibility_score: A self-reported measure of participants’ tendency to be gullible.
  • Big Five Personality Traits: A measure including extraversion, openness, agreeableness, conscientiousness, and emotional stability.
  • interest_behavior: A measure of participants’ interest in the topic of prediction (personal investment behavior).
  • familiarity: A measure of participants’ familiarity with the prediction sources (AI, Astrology, Personality).
  • age
  • gender: A categorical variable (Female, Male, Other)
  • education: A categorical variable (Bachelor, Master, Doctorate, Professional degree, Associate, Some college, High school, Less than high school)
7.5.2 Multiple linear regression
To test the hypothesis that individuals who are more likely to believe in astrology and personality-based predictions are also more likely to believe in AI predictions (H1), a multiple linear regression model was first applied to the data in wide format. The dependent variable was ai_overall_score, and the main predictors were zodiac_overall_score and personality_overall_score, with the rest of the variables mentioned in Section 7.5.1 included as control variables. The model was fitted using the lm function in R.
[…]
The following model diagnostics were performed to ensure the validity of the model:
  • Linearity: Residual plots for the linear regression model indicated no clear patterns, suggesting linearity.
  • Homoscedasticity: The residual vs. fitted values plot for the linear regression model showed some heteroscedasticity, which was addressed in the mixed-effects model.
  • Normality of Residuals: The Q-Q plot indicated that the residuals were approximately normally distributed.
  • Multicollinearity: Variance Inflation Factors (VIFs) were calculated for the linear regression model, indicating that multicollinearity was not a concern.
  • Independence of Residuals: The Durbin-Watson test for the linear regression model showed no significant autocorrelation.
7.5.3 Mixed effects model
To examine the main effects and moderating effects of psychological and cognitive factors (H2, H3), the data was transformed to a long format, creating a three-level hierarchical structure. Each subject engaged in three types of conditions (prophecy_source) and evaluated across four subscales (subscale). The outcome variable was the subscale_score.
The mixed effects model was specified as follows. The fixed effects were defined as:
[…]
For the random effects, random intercepts were included to account for variability in baseline levels between subjects (qualtrics_code), and random slopes were included to account for variability in how subjects respond to different conditions (prophecy_source):
[…]
Additionally, the model incorporated a correlation structure and variance weights to further account for the hierarchical structure and heteroscedasticity in the data:
[…]
We centered the continuous predictor variables to reduce multicollinearity and improve the stability and interpretability of the coefficient estimates. Centering involved subtracting the mean of the variable from each individual value. This transforms the variable to have a mean of zero while preserving its variance and distribution.
After standardizing the predictors, we observed that four previously non-significant main and interaction terms became significant, including the main effect of Personalization subscale, the main effect Astrology prediction source, and interaction terms between subscale (Personalization, Reliability) and Astrology prophecy source. This change is attributed to the reduction in multicollinearity, resulting in more precise coefficient estimates. The model fit statistics (AIC, BIC, log-likelihood) indicated an overall improvement in model fit after standardization (AIC_before = 7949.6, AIC_after = 7693.4; BIC_before = 9129.2, BIC_after = 8873.0; logLik_before = -3774.8, -3774.7_after = -3646.7).
The model was implemented using the lme function from the nlme package in R. The analysis controlled for missing values through listwise deletion. Optimization settings were adjusted to ensure convergence, with maximum iterations and evaluations set to 1000.
Several model diagnostics were performed to ensure the validity of the model:
  • Linearity: A residuals vs fitted values plot using standardized residuals indicated no clear patterns, suggesting linearity.
  • Homoscedasticity: A residuals vs fitted values plot showed a random spread around the horizontal line at zero, confirming that heteroscedasticity was successfully addressed using the variance function (varIdent). One limitation to this diagnostics was the less continuous nature of the outcome variable (subscale_score), which was multilevel with 13 levels (value of 1 to 7 with an increment of 0.5), which caused the residual plot to exhibit a slight diagonal pattern.
  • Normality of Residuals: Normality of residuals was assessed using a Q-Q plot, using standardized residuals to account for both fixed and random effects. While there were slight deviations at the extremes, they were not severe and thus considered as normal deviations from real-world data.
  • Multicollinearity: To address potential structural multicollinearity, we centered the continuous predictor variables. This process reduced the Variance Inflation Factor (VIF) values for all fixed effects to below 5, indicating that multicollinearity was reduced to an acceptable level. The model diagnostics, including standard errors and confidence intervals, confirmed the stability and reliability of the coefficient estimates.
  • Independence of Residuals: We applied the Durbin-Watson test within each group at the lowest level (subscale) to check for autocorrelation. The test results indicated significant positive autocorrelation for each group. To address this, we specified an autoregressive correlation structure (corAR1), which slightly mitigated the autocorrelation. However, model fit statistics (AIC, BIC, and log-likelihood tests) suggested that using an unstructured symmetric correlation structure (corSymm) provided a better fit than the autoregressive structure. Therefore, we chose the unstructured symmetric correlation as the final model. A comparison of the p-values and estimates between models with corAR1 and corSymm structures revealed no notable differences.
    While the corSymm structure improved overall model fit, it did not fully address the positive autocorrelation in the lowest level residuals. This unresolved autocorrelation may pose a limitation to the interpretation of our results, as it could affect the validity of statistical inferences. Future studies could explore alternative approaches or additional data collection to better address this issue.
  • Normality of Random Effects: Q-Q plots were used to check the normality of random effects. The Q-Q plots of random effects showed that the points followed the reference line closely, suggesting normality. There were slight deviations at the tails, suggesting minor departures from normality, but these deviations were not deemed significant enough to invalidate the model assumptions.
  • Random Effects Evaluation: Variance components and Intraclass Correlation Coefficient (ICC) were examined to evaluate the random effects structure. The ICC values (Adjusted ICC: 0.939, Unadjusted ICC: 0.598) indicated that a significant portion of the variance was attributable to the random effects. To assess the contribution of random effects to the overall model variability, we examined the variance components of the mixed effects model. The model included random intercepts and slopes for prophecy_source at the level of qualtrics_code. The estimated variance components are presented in Table 4.
Random Effect Variance SD Corr w Intercept Corr w Astrology
Intercept 1.560 1.249 - -
prophecy_sourceAstrology 1.211 1.100 -0.260 -
prophecy_sourcePersonality 1.002 1.001 -0.221 0.293
Residual 0.121 0.348 - -
Table 4: Random Effects Variance, Standard Deviation, and Correlations
Acknowledgments
We thank Jinjie Liu, Data Science Specialist at the Institute for Quantitative Social Science at Harvard University for sharing her valuable perspective on the analysis.